OpenStudy (anonymous):

If y=4(1.6)^x , what is the approximate value of x when y = 12? Answer a. 2.1 b. 1.9 c. 2.5 d. 2.3

OpenStudy (callisto):

y=4(1.6)^x and y =12 12 = 4(1.6)^x 3=1.6^x xlog(1.6) = log3 x = log3 / log1.6 =?

OpenStudy (anonymous):

b?

OpenStudy (callisto):

Nope :|

OpenStudy (anonymous):

thats what i got when i typed it in what u get?

OpenStudy (callisto):

x=(log3) / (log1.6)

OpenStudy (anonymous):

just plug the answers they gave you in for x and when you get 12 as your answer, that is the x....does that make sense

OpenStudy (anonymous):

12=4(1.6)^x a. 2.1=x b. 1.9=x c. 2.5=x d. 2.3=x

OpenStudy (anonymous):

so C

OpenStudy (anonymous):

so which is it one of your answers is c and the others is d

OpenStudy (callisto):

Actually, instead of putting the value into the equation provided, you can solve it :|

OpenStudy (anonymous):

either way, that's what i do on tests...just so i make sure i get the right answer...ha sorry if i made more work for you :/